OpenStudy (anonymous):

Multiply 14x/a^4 * 2a^5 x/7x^3 simplify answer as much as possible

OpenStudy (anonymous):

@Euler271

OpenStudy (anonymous):

\[x^a*x^b = x^{a+b}\]\[\frac{ x^a }{ x^b } = x^{a-b}\] \[\frac{ 14x*2a^5x }{ a^4*7x^3 } = \frac{ 28 }{ 7 }*\frac{ x^2 }{ x^3 }\frac{ a^5 }{ a^4 } = \frac{ 4a }{ x }\]

OpenStudy (anonymous):

\[x^{-1} = \frac{ 1 }{ x }\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i cant typee =

OpenStudy (anonymous):

is the answer 4a/x ?

OpenStudy (anonymous):

@Euler271

OpenStudy (anonymous):

yes